### v1

# HateDay: Insights from a Global Hate Speech Dataset Representative of a Day on Twitter

Abstract
To tackle the global challenge of online hate speech, a large body of research has developed detection models to flag hate speech in the sea of online content. Yet, due to systematic biases in evaluation datasets, detection performance in real-world settings remains unclear, let alone across geographies. To address this issue, we introduce HateDay, the first global hate speech dataset representative of social media settings, randomly sampled from all tweets posted on September 21, 2022 for eight languages and four English-speaking countries. Using HateDay, we show how the prevalence and composition of hate speech varies across languages and countries. We also find that evaluation on academic hate speech datasets overestimates real-world detection performance, which we find is very low, especially for non-European languages. We identify several factors explaining poor performance, including models’ inability to distinguish between hate and offensive speech, and the misalignment between academic target focus and real-world target prevalence. We finally argue that such low performance renders hate speech moderation with public detection models unfeasible, even in a human-in-the-loop setting which we find is prohibitively costly. Overall, we emphasize the need to evaluate future detection models from academia and platforms in real-world settings to address this global challenge.
Content warning: This article contains illustrative examples of hateful content.
1 Introduction
Social media users often encounter hate speech in their feeds UNESCO (2023), fueling fears of its impacts on offline violence Müller and Schwarz (2021). To counter this phenomenon, a mature body of research has developed automatic hate speech detection models Vidgen and Derczynski (2020). Yet, past work in this space suffers from three major shortcomings.
Figure 1: HateDay consists of twelve annotated representative sets (N=20K each) randomly sampled from all tweets posted on September 21, 2022 in eight languages (Arabic, English, French, German, Indonesian, Portuguese, Spanish, and Turkish) and originating from four countries where English is the main language on Twitter (United States, India, Nigeria, Kenya).
First, the performance of hate speech detection systems in real-world settings remains largely unclear. On the one hand, detection models used in industry are not public and detection performance metrics are usually missing from platform transparency reports or sometimes misleading Giansiracusa (2021). On the other hand, evaluation of hate speech models in academic research has been done on biased datasets whose characteristics greatly differ from real-world settings, from class imbalance to topic distribution Nejadgholi and Kiritchenko (2020), raising concerns about an overestimation of the reported detection performance Arango et al. (2019); Wiegand et al. (2019).
The second major limitation of academic research on automatic detection of hate speech lies in its limited linguistic scope, with most resources developed for English Poletto et al. (2021); Tonneau et al. (2024a). While such unequal resource allocation is likely to result in unequal performance across languages, the use of customized datasets for evaluation renders cross-lingual performance comparisons unfeasible.
Finally, the focus on languages in the development of hate speech detection systems may hide performance differences between countries that share a common language Ghosh et al. (2021). While recent work has focused on developing datasets grounded in geographical location rather than language Maronikolakis et al. (2022); Castillo-lópez et al. (2023), cross-country performance differences remain largely unclear. In this context, it is necessary to develop multilingual evaluation sets that are representative of real-world settings and that provide a comparable setting across languages and countries with a common language.
In this article, we aim to address these three shortcomings and provide an evaluation of state-of-the-art hate speech detection performance on real-world social media data across languages and countries. For this purpose, we introduce HateDay, a global Twitter dataset consisting of 240,000 tweets sampled from all tweets posted on September 21, 2022 and annotated for hate speech detection. Specifically, we draw random samples of 20,000 tweets each for eight popular languages on Twitter: Arabic, English, French, German, Indonesian, Portuguese, Spanish, and Turkish. In addition, we draw random samples of 20,000 tweets each for four countries whose major language on Twitter is English: the United States, India, Nigeria, and Kenya.
We show that the prevalence and composition of hate varies across languages and countries. We then use HateDay to evaluate public hate speech detection models across languages and countries. We find that hate speech detection performance is largely overestimated by evaluation on academic hate speech data from prior work and is very low in real-world social media settings, particularly for non-European languages. We identify several factors explaining such poor model performance including the inability of models to distinguish between hate and offensive speech, and the misalignment between academic target coverage and real-world target prevalence. Finally, in light of our performance results, we estimate the feasibility of hate speech moderation using publicly available models. We argue that automatic moderation alone is too error-prone to be desirable and show that human-in-the-loop moderation, whereby models flag potentially hateful content for human verification, proves prohibitively costly across languages and countries, despite its potential for improved accuracy. In sum, our contributions are:
  1. 1.
    HateDay, the first global hate speech dataset representative of social media settings [^1] Available at https://huggingface.co/datasets/manueltonneau/hateday, composed of 240,000 annotated tweets covering eight languages and four countries.
  2. 2.
    a comparison of hate speech prevalence and composition across languages and countries
  3. 3.
    an estimation of cross-lingual and geographic hate speech real-world detection performance, including a detailed analysis of pitfalls of current detection models
  4. 4.
    an estimation of the feasibility of hate speech moderation using publicly available models
Figure 2: Prevalence of harmful content (left) and targets of hate speech (right) in HateDay. The target category “Other” contains rare target labels such as “Caste”, “Age”, “Occupation”, “Disability” and “Social Class”.
2 The HateDay Dataset
2.1 Data Collection
As a basis for creating HateDay, we use the TwitterDay dataset Pfeffer et al. (2023), which contains all Twitter posts within a 24-hour period starting on September 21, 2022. This amounts to approximately 375 million tweets posted by 40 million users platform-wide.
We filter TwitterDay both at the language and country levels to enable a more structured analysis. At the language level, we focus on the eight most popular languages in the TwitterDay dataset for which there exist hate speech detection resources (details in §A.1), namely Arabic, English, French, German, Indonesian, Portuguese, Spanish and Turkish. To assess differences between countries with a common language, we also filter at the country-level, focusing on four countries for which English is the main language on Twitter, namely the United States, India, Nigeria and Kenya. For locating users in each country, we use the Google Geocoding API to infer user country location based on their self-provided profile location Hecht et al. (2011). We drop retweets to focus on original content and randomly sample 20,000 tweets for each of the eight retained languages and each of the four countries, ensuring that each language-specific and country-specific sample is representative of real-world Twitter settings in that particular language or country. The HateDay dataset corresponds to the combination of all the annotated random samples for each language and country, for a total of 240,000 annotated tweets.
2.2 Data Annotation
We recruit a team of 36 annotators, three per language or country. For languages that are spoken widely around the globe, such as English, Arabic or French, we maximize the diversity of annotator origins where possible (details in §B.1).
We follow a prescriptive approach to manage annotator subjectivity in our annotation task Rottger et al. (2022) by instructing annotators to strictly adhere to extensive annotation guidelines that describe our taxonomy of hate speech (provided in §B.2). Following prior work Davidson et al. (2017); Mathew et al. (2021), the annotation consists of assigning each tweet to one of three classes: (i) hateful, if it promotes violence or contains a direct attack, abuse or threat against an individual or a group based on the perceived belonging of a certain characteristic (e.g., gender, race), (ii) offensive, if it contains any form of non-acceptable language that is not hateful, including insults, threats, and posts containing profane language or swear words Zampieri et al. (2019), or (iii) neutral if it is neither hateful nor offensive. For tweets labeled as hateful, we also ask annotators to specify the groups being targeted. We provide examples of tweets for each class in Table 1.
Class Target Category Examples Origin
Politics Greens are vermin German
National Origin Venezuelans are useless Spanish
Hateful Gender I hate women Portuguese
Religion Hunt these muslim b*stards India
Sexual Orientation F*ck these f*ggots French
Offensive n/a Stop spewing rubbish, fool. Nigeria
Neutral n/a Reunited by the mercy of God Turkish
Table 1: Examples of tweets for each class and main target categories. Offensive tweets, by definition, have no target. English translations are displayed in italic.
For each language and country, we conduct a pilot annotation to train annotators on the task. Specifically, we have each annotator label 100 posts sampled from a collection of all annotated hate speech datasets Tonneau et al. (2024a) open-sourced on Hugging Face, a major open source platform. We repeat this task until the inter-annotator agreement, measured by Krippendorff’s $\alpha$, reaches 0.7. After the pilot, we task each annotator to label the random sample of 20,000 tweets in their respective language or country. Each tweet is labeled by three annotators, and the final label is determined by majority vote. To minimize the amount of hate that is wrongly categorized as not hate, all tweets without a majority vote or with at least one hateful label out of three are re-examined by all three annotators and a task expert to derive their final label. Across all languages and countries, the three annotators agreed on 93.4% of all labeled tweets, two out of three agreed in 6.4% of cases, and all three of them disagreed in 0.2% of cases.
2.3 Descriptive Statistics
Prevalence of harmful content
We find that the prevalence of hate speech on Twitter on the day of analysis is very low, representing less than 2% of all posts for all considered languages and countries and with an average prevalence of 0.7% (Figure 2, left). In contrast, we observe that offensive content is much more prevalent than hate speech across all languages and countries, ranging from 2.5 times more prevalent for French to 19 times more prevalent for Portuguese. We also observe large differences in the prevalence of hate speech between languages and countries. For instance, 1.0% of German tweets are hateful whereas only 0.3% of Indonesian tweets are hateful. We observe similar gaps at the country level, with the share of hateful posts being much higher in India (1.7%) than in Kenya (0.2%) or Nigeria (0.3%).
Main targets of hate speech
We find that the most common targets of hate in HateDay are political as well as racial and gender groups (Figure 2, right). We also find notable differences across languages and countries. For instance, political hate speech is prevalent in English, German, Indonesian, Portuguese, and Spanish, as well as in the U.S. national context, whereas it is less present in Turkish and Nigerian tweets, and not present at all in Arabic tweets. Also, religious hate speech, mostly in the form of Islamophobia, dominates in India whereas it is less prominent elsewhere and completely absent from the sample in the Kenyan case. Finally, we note that some forms of hate speech are unique to specific contexts, such as casteism in India.
3 Experimental Setup
Across our experiments, we leverage the language and country-level representativeness of the HateDay dataset to evaluate public hate speech detection models in real-world social media settings.
3.1 Models
We evaluate models that are either trained for the task of hate speech detection (supervised learning) or prompted for this task (zero-shot learning).
Supervised learning
For each language and country, we identify and evaluate all hate speech detection models that are open-sourced on Hugging Face, and trained using supervised learning. In cases where there are more than five models, we limit our analysis to the five most downloaded models on Hugging Face at the time of analysis (August 2024). We provide the full list of open-source models that we evaluated in §D.1. Additionally, we include the Perspective API Lees et al. (2022), a widely used toxic language detection system also based on supervised learning.
Zero-shot learning
We use Aya23 8B Aryabumi et al. (2024) and Llama3.1 8B Dubey et al. (2024) for zero-shot learning. We do so because Aya is designed to be multilingual, and Llama3.1 is one of the most capable open-source decoder-only models available at the time of the analysis. We use smaller architectures due to computing constraints and provide the prompt for zero-shot learning in §D.2.
3.2 Evaluation
We evaluate models on the HateDay (HD) datasets to estimate real-world detection performance. Additionally, we compare the latter with the performance on two traditional evaluation methods, namely academic hate speech datasets and functional tests.
Academic hate speech datasets (AD)
We measure performance on academic hate speech datasets to understand how results from past work generalize to our more realistic HateDay datasets. We rely on open-sourced hate speech supersets combining all existing hate speech datasets for all eight languages of interest Tonneau et al. (2024a). Given that English hate speech datasets mostly originate from the US Tonneau et al. (2024a), we use English hate speech datasets both for evaluating English-centered and US-centered models. In the absence of supersets for India, Nigeria and Kenya, we survey all existing datasets for each country and combine them to build the supersets (details in §C.2). We restrict the evaluation to a 10% random sample of all academic hate speech datasets for each language and country to limit the computing resources required for inference (details in §C.1).
Functional tests (HC)
We measure performance on functional tests to estimate models’ ability to handle known challenges in hate speech detection. We use HateCheck Röttger et al. (2021, 2022), a suite of functional tests for hate speech detection models, which covers six of eight languages in HateDay, missing Indonesian and Turkish.
We evaluate model performance on each dataset using average precision, which corresponds to the area under the precision-recall curve, and is well suited when class imbalance is high, as is the case in HateDay.
4 Results
4.1 Detection Performance
Model Type Arabic English French German Indonesian Portuguese Spanish Turkish
AD HC HD AD HC HD AD HC HD AD HC HD AD HC HD AD HC HD AD HC HD AD HC HD
Llama 3.1 10 76.4 1.2 38.8 88.5 4.6 33.5 84.9 7.3 18.9 84.2 5.8 62.8 - 2.5 15.5 84.8 2.2 39.4 84.9 7.2 39.3 - 2.9
Aya 23 11.5 77.6 2 36.1 87.4 3.5 37.1 84 5.3 21.6 86 5.9 61.6 - 2.6 15.9 86.6 2.2 40.7 85.3 7.6 44.6 - 4.5
Best OS 25.8 76.9 5.8 38.2 82.2 9 35.7 89.9 16 64.3 97.2 19.3 90.2 - 3.5 18.7 85.8 3.1 54.9 82.3 11.6 32.1 - 1.9
Perspective 18.7 89.6 10.2 58.9 95.1 10.1 40.1 96.9 37.7 53.6 96 18.8 65.5 - 11.1 30.9 94.6 14.5 50.9 96 34.1 $31.4^{*}$ $\textbf{6.5}^{*}$
Table 2: Model performance across languages and evaluation sets, as measured by average precision (%). We report performance on three evaluation sets: academic hate speech datasets (AD) combined for a given language, HateCheck functional tests (HC) and HateDay (HD). HC does not cover Indonesian and Turkish. Asterisks indicate that the language of interest is not supported by the Perspective API.
We report the average precision of each model across the three aforementioned datasets, across languages (Table 2) and across countries (Table 3). We also provide results on individual open-source model performance in §E.2.
Performance across datasets
Our most striking finding is that across the wide range of models, languages and countries considered in this study, the average precision on our representative HateDay data (HD) is just 9.4% on average, compared to 40% on the sample of public hate speech datasets (AD) and 87.2% on the functional tests (HC).
Performance across model types
We find that supervised learning consistently outperforms zero-shot learning across almost all combinations of language, country, and evaluation sets. Indeed, the best open-source model on HateDay has an average precision of 36.0% on average across all combinations, whereas Llama3.1 8B, which performs on par with Aya23 8B, scores 28.4% on average. Additionally, despite being originally developed to detect toxic rather than hateful content, we observe that the Perspective API (39.1% average precision on average) outperforms open-source hate speech detection models in all languages but German. Still, the dominance of Perspective is more limited at the country-level where it performs best in the US and Indian contexts but not in the Nigerian context, and only in the Kenyan context because there are no supervised models to compare to.
Model Type United States India Nigeria Kenya
AD HD AD HD AD HD AD HD
Llama 3.1 38.8 4.9 50.9 13.4 32 2.6 24.3 1.5
Aya 23 36.1 4.7 52.7 10.4 30.7 1.6 23.9 0.8
Best OS 38.2 9.8 54.3 7.8 65.7 30.9 - -
Perspective 58.9 12.3 61.9 42.9 44.1 8.6 31.6 9.1
Table 3: Model performance across countries and evaluation sets, as measured by average precision (%). There is no open-source model specifically for Kenya.
Cross-lingual and geographic gaps
We observe substantial performance differences in the real-world setting of HateDay across languages and countries. At the language-level, we find that the best average precision for European languages (23.1% on average), such as French, Spanish or German, is substantially higher than for non-European languages (9.3%), such as Arabic, Indonesian and Turkish. At the country-level, the average precision is by far highest on Indian tweets (42.9%) followed by Nigeria (30.9%), whereas the performance is rather low in both the U.S. (12.3%). and Kenyan (9.1%) contexts.
4.2 Performance Analysis
In the following, we conduct a more in-depth analysis into potential explanations for the poor hate speech detection performance as well as the cross-geographic performance gaps that we observe in real-world settings in §4.1.
Low precision and recall
As the average precision corresponds to the area under the precision-recall curve, the same average precision may correspond to very different patterns in precision and recall. We therefore inspect the precision-recall curves of the best-performing models on HateDay for each language and country. We find that, while precision or recall may be higher depending on the context, both values remain very low and there are no situations where both values are above 50%, with the highest F1-score at 47% for India (Figure 5 in the Appendix). This highlights the role of both false positives and negatives in the observed low performance. Next, we study the composition of these two types of errors.
Figure 3: Share of hateful, offensive and neutral content in the top 5% scored tweets (N=1,000) in HateDay for each language and country. We use the score distribution of the best performing model on HateDay (Tables 2 and 3). The x-axis corresponds to the descending rank and each bar shows the distribution of content in a window of 50 tweets.
Offensive false positives
We find that offensive content constitutes a substantial share of the top of the score distribution — 32.3% in the top 50 (0.25%) scored tweets on average — for each language and country (Figure 3), thereby hurting precision. For Arabic, English, Portuguese and the U.S. and Kenyan national contexts, offensive content even appears more frequently than hate speech at the top of the score distribution. This problem is further aggravated by the fact that offensive content is more prevalent than hate speech on the day of the analysis (Figure 2). Indeed, we observe that as the prevalence ratio of offensive content to hate speech increases, the share of offensive content also gradually increases (Figure 7 in the Appendix).
Qualitative analysis of false positives
We further examine top-scored neutral tweets to identify other prominent features of false positives. We find that many such tweets contain mentions of hate speech (e.g., “wasn’t it the political scientist saying that Ukrainians are “subhumans”?”; “for them, women are like joints, they smoke them until they get high”). We also find several cases of statements that are not hateful but that would become hateful with a few changes of letters (e.g., “emigration is a catastrophy”; “this was noggers, man [^2] Portuguese original is: “isso foi noggers mano.”. “Noggers” means “disappointing” in Portuguese gamer slang.). Finally, we find that a substantial share of such tweets may be hateful but lack the context to fully determine their hatefulness. Such tweets often are replies and contain target community names (e.g., “@USER No, illegal migrants, they say it themselves”; “@USER We know you are gay”).
Analysis of false negatives
Next, we analyze false negatives, defined as hateful examples that do not appear in the top 1% of the best model’s score distribution for each language and country. We first examine the targets of these tweets to determine how their representation in false negatives compares to their overall prevalence in hate speech for each specific language or country. We find that politics and gender are overrepresented in false negatives, respectively in 10/12 and 7/12 of the studied language and country-level cases. In parallel, religion and race are under-represented, respectively in 10/12 and 8/12 of cases. We also look at ambiguity, considering that a hateful example is ambiguous if not all three annotators labeled it as hateful. The share of ambiguous content is significantly higher in false negatives compared to the whole set of hateful examples across languages and countries (t(11) = 3.76; p-value < 0.05).
Target academic focus and prevalence
Motivated by the overrepresentation of certain targets in false negatives, we inspect the role of target-level academic focus, as a proxy for the distribution of hate types seen by the model during training, on performance. We first examine differences at the target level between the attention each target group has received in academic work and its prevalence in HateDay across languages and countries. We use data from a recent survey which documents the share of all hate speech datasets by target category and language Yu et al. (2024). In the absence of data at the country level, we categorize all surveyed datasets in terms of the target focus for each country of interest (details in §C.2.2). For each language and country, we provide a comparison between the target-level shares of (i) all hate speech datasets and of (ii) all hate speech in HateDay in Figure 6 in the Appendix. Our most notable finding is that political hate speech is much more prevalent in HateDay in two-thirds of cases than it is studied in academic work on hate speech. To a lesser extent, we also find that gender-based hate is often more prevalent in HateDay than is studied while religion-based hate is less prevalent than is studied in existing hate speech datasets.
Finally, we explore the impact of the alignment between the target focus of academic work and the target prevalence in real-world social media data on performance. We estimate such alignment for each language and country by computing the cosine similarity between the vectors containing the share of all hate speech datasets by target focus and the share of all hate speech in HateDay by target. We find a strong and significant positive correlation between this alignment and detection performance (Pearson’s r: 0.76, p-value: 0.003). This indicates that the better hate speech detection resources reflect hate target coverage in real-world social media data for a given language or country, the higher the detection performance. In contrast, we do not find a significant correlation between the amount of annotated datapoints in the aforementioned supersets and performance on HateDay (Pearson’s r: -0.47, p-value: 0.127). This indicates that the amount of detection resources plays a smaller role in cross-lingual and geographic performance differences.
4.3 Feasibility of Hate Speech Moderation
Figure 4: Cost-recall tradeoff in human-in-the-loop moderation. Share of all HateDay tweets flagged as hateful and reviewed by moderators (%) versus share of all hate in HateDay moderated (%). We use the best performing model on HateDay for each language and country (Tables 2 and 3). The dashed line indicates performance of a model that flags tweets as hateful at random.
Most social media platforms prohibit hate speech Singhal et al. (2023) and have done so since their inception Gillespie (2018). In light of our performance results, we investigate the feasibility of hate speech moderation using publicly available hate speech detection models. Specifically, we consider human-in-the-loop moderation, which is an approach used by large social media platforms Avadhanula et al. (2022) whereby content likely to be hateful is flagged by detection models for review by human moderators. We study the trade-off between the amount of human reviewing required and the corresponding share of all hate successfully moderated (Figure 4, details in E.3).
We find that successfully moderating a large share of all hateful content (>80%) using the best publicly available detection models would require human review of least 10% of all tweets posted in a given day. For such a review workload, the total share of hate successfully moderated varies between 70% and 90% for most languages or countries, apart from Turkish where this share is only 40%. 10% of all daily tweets in a given language represents a massive amount of posts, ranging from 19.3 thousand for Kenya to 5.1 million for English. A smaller share of human review would only allow to moderate a small share of all hate: when flagging and reviewing 1% of all tweets posted in a given day, only 20-40% of all hate are moderated for most languages and countries.
5 Discussion
Hate speech prevalence
We find that the proportion of hate speech relative to all social media content is very low, which is in line with past work Gagliardone et al. (2016); Mondal et al. (2017); Park et al. (2022). This contrasts with the high share of social media users who report that they often encounter hate speech in their feeds Vidgen et al. (2019); UNESCO (2023) and points to the potential role of recommendation algorithms in amplifying harmful content Milli et al. (2023).
Performance overestimation
We find that traditional methods largely overestimate performance in randomly drawn samples of real-world social media data, which is very low, especially for non-European languages. While past work has discussed the risk of such overestimation Arango et al. (2019), due to the biased nature of hate speech datasets Wiegand et al. (2019); Nejadgholi and Kiritchenko (2020), and quantified such risk in the Nigerian Twitter context Tonneau et al. (2024b), we provide evidence of low and overestimated performance across many more linguistic and national contexts. Moreover, our results highlight the need to evaluate NLP tools in realistic settings, especially where human harm might arise. This applies more broadly to the detection of harmful content on social media, such as misinformation Thorne et al. (2018), as well as the evaluation of the bias and safety of generative large language models Ibrahim et al. (2024); Lum et al. (2024); Röttger et al. (2024), both of which have relied so far on non-representative benchmarks for evaluation.
Dominance of Perspective API
We find that the Perspective API often outperforms other models in hate speech detection across languages, contradicting existing evaluations on traditional hate speech datasets showing that Perspective is outperformed by academic hate speech classifiers Wich et al. (2022); OFCOM (2024). We attribute such dominance to the fact that Perspective is optimized for real-world performance and generalizability to unseen data Lees et al. (2022), contrary to traditional hate speech detection models developed in academia. We also note that the Perspective API is periodically updated and that performance may have improved since the evaluations conducted in past work Pozzobon et al. (2023). Yet, we also observe that Perspective’s dominance does not hold at the country level. This highlights the limits of the language-level approach in developing Perspective and more generally hate speech classifiers Tonneau et al. (2024a), which may not be tailored to national contexts where non-US English and code-mixing is prevalent.
The problem of offensive language
We find that a major reason for the poor performance of hate speech classifiers on HateDay is the high prevalence of offensive content at the top of the hate score distribution, which reduces precision. While offensive content differs from hate speech because it does not target based on the perceived belonging of a particular characteristic, it shares semantic similarities such as the use of swear words and profanities. These similarities likely explain the high hate scores given by classifiers to offensive content. While prior work initially documented this issue Davidson et al. (2017), our findings demonstrate its heightened relevance in social media settings, aggravated by the substantially higher prevalence of offensive content compared to hate speech.
Unseen targets of hate
Our findings reveal a misalignment between the targets of hate speech studied in academic work and their real-world prevalence in HateDay. Notably, political-related hate is understudied in academic work compared to its prevalence in HateDay, expanding on past work that points to a low representation of certain targets in hate speech detection resources Yu et al. (2024). We also find that the alignment between academic target coverage and real-world target prevalence has a significant positive correlation with detection performance in HateDay, contrary to the raw amount of detection resources. Improving such alignment, particularly for underrepresented hate types such as political-based hate, is therefore crucial for enhancing real-world detection performance. In the process, political criticism and hate speech must be clearly differentiated to avoid accusations of politically biased moderation Vogels et al. (2020), and hate speech laws being misused for political censorship Strossen (2018).
Other challenges for hate speech detection
Our analysis identifies additional challenges for hate speech detection partly identified in past work: First, the difficulty of distinguishing between the use and the mention of hate speech, for instance as part of counter speech Röttger et al. (2022); Gligoric et al. (2024); Jin et al. (2024). Second, the inability of classifiers to capture subtle differences in wording that change a statement from hateful to non-hateful. Third, the lack of context in many posts, especially replies, poses challenges to classification, and calls for more contextualized hate speech detection Pérez et al. (2023).
Implications for hate speech moderation
Our results indicate that hate speech moderation using public hate speech detection models is ineffective at curbing online hate speech at a limited cost. Indeed, low real-world detection performance renders fully automated moderation ineffective and potentially harmful, as it fails to protect users from hate speech while likely censoring non-hateful content, such as counter-speech, confirming concerns from past theoretical work Gorwa et al. (2020). We also show that human-in-the-loop moderation based on public detection models can only effectively moderate a large share of hate speech across our languages and countries of interest at a prohibitively high annotation cost, extending past results for the Nigerian Twitter context Tonneau et al. (2024b) to more languages and countries. Beyond financial costs, the required scale of human review also raises concerns about potential reviewer harm from repeated exposure to harmful content Kirk et al. (2022). Our results also support the claim that detection alone will not solve the hate speech problem Parker and Ruths (2023) and call for complementary solutions to this problem. These include preventive approaches which aim to reshape speech norms and have proven effective to curb hate speech, for instance by prompting users to reconsider harmful posts Katsaros et al. (2022), or by confronting hate speech with counter speech Munger (2017); Hangartner et al. (2021). Finally, we note that our results do not imply that hate speech moderation itself is ineffective, because social media platforms use their own private models for moderation and have access to more data which may improve performance (see Limitations). Yet, without access to real-world performance metrics for platform models, the effectiveness of hate speech moderation on social media remains uncertain.
6 Conclusion
In this article, we introduced HateDay, the first global hate speech dataset representative of real-world social media settings. We found that evaluations based on previous academic datasets overestimate real-world performance, which is very low. We further discussed the implications of our results for hate speech moderation, namely that it is unfeasible using public detection models. In light of our results, future work is needed to improve hate speech detection performance. In the process, we encourage researchers to ground their evaluation in the real-world settings in which their models may eventually be used. We also encourage platforms to be more transparent on how their detection models perform in social media contexts, in order to better understand the effectiveness of hate speech moderation. We hope that our dataset and experimental results will stimulate work in both of these directions to help solve this important problem.
Limitations
Dataset
Low number of positives: The random samples in HateDay used to evaluate hate speech detection in real-world settings contains a low number of hateful examples, ranging from 31 for Kenya to 430 for India. This low number is linked to the very low prevalence of hate speech in our dataset as well as our budget constraint which impeded us from further expanding the annotation effort. While statistically significant, we acknowledge that our performance results on HateDay are necessarily more uncertain, as illustrated by the larger confidence intervals in Tables 5 and 6.
Limited generalizability to other platforms, timeframes, and linguistic domains: The entirety of our dataset was sampled from one single social media platform for a very bounded timeframe, namely 24 hours. This limits the generalizability of our performance results to data from other social media platforms and covering other timespans.
Online hate speech is multimodal: Our work focuses only on text-based hate to limit annotation costs, but we acknowledge that online hate speech is multimodal and that a non-trivial share of this phenomenon, which our analysis necessarily misses, is expressed through images, sounds, or videos Hee et al. (2024).
Limits to representativeness: The tools used to stratify the TwitterDay dataset into language or country-specific sets, namely language detection and user location inference, are imperfect Hecht et al. (2011); Graham et al. (2014); Jurgens et al. (2017). This implies that the representativeness of HateDay may not be perfect, with language or country sets containing some posts in other languages or from other countries. We acknowledge this limitation but argue that HateDay remains as close as a set of representative samples of Twitter as can be with existing stratification tools.
Moderation prior to data collection: Our analysis assumes that the hateful content in HateDay is representative of all hateful content posted on Twitter on the day of analysis. However, we recognize that some instances of hate speech may have been moderated by Twitter before the data was collected, making our estimates a lower bound. Nonetheless, since the posts were collected 10 minutes after they were posted Pfeffer et al. (2023), we believe that the enforcement of moderation in such a short timeframe is likely to be minimal.
Moderation analysis
We acknowledge that our analysis on the feasibility of moderation is limited in the sense that we operate with publicly available resources while platforms also rely on private models and data which may improve moderation performance for a given cost.
Private models While we evaluate publicly available detection models, we are aware that platforms have developed their own detection models, which are kept private and may perform better.
Private data Platforms also rely on several data sources to develop their moderation models, such as the user history, user-graph, conversational context, which we do not have access to. While we consider only public hate speech detection algorithms as a way to flag hate speech in the flow of social media content, we are also aware that platforms use other flagging mechanisms based on private data, such as user reporting or banned word lists, which may increase the share of all hate moderated for a given annotation cost. Finally, the metric platforms aim to reduce is the view count of harmful content such as hate speech, which is arguably more relevant than the sheer prevalence of it from a harm perspective but is also private for the considered timeframe.
Ethical Considerations
Annotator wellbeing
Annotators were provided with clear information regarding the nature of the annotation task before they began their work. They were made aware that they could stop the task at any time if necessary.
Data privacy
To protect the identity of hateful users and their victims, we anonymize all tweets in our dataset upon release, replacing all user names by a fixed token @USER.
Intended use
The HateDay dataset will be shared publicly upon publication and its intended use is for research purposes only.
Acknowledgements
We thank the annotators for their excellent research assistance. We also thank Dhruti Sanghavi for outstanding visual design work. The study was supported by funding from the United Kingdom’s Foreign Commonwealth and Development Office (FCDO), the World Bank’s Research Support Budget, and the Bill and Melinda Gates Foundation. This work was also supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. The findings, interpretations, and conclusions expressed in this paper are entirely those of the authors. They do not necessarily represent the views of the International Bank for Reconstruction and Development/World Bank and its affiliated organizations or those of the Executive Directors of the World Bank or the governments they represent.
Appendix A Twitter Day
A.1 Language distribution
We provide the language share of all posts in TwitterDay Pfeffer et al. (2023) in Table 4. We retain the languages for which there was one at least one academic hate speech dataset on https://hatespeechdata.com/ at the time of the analysis (August 2024).
Language Share (%)
English 27.6
Japanese 20.9
Spanish 6.7
Arabic 6.6
Portuguese 5.4
Indonesian 2.9
Korean 2.5
Turkish 2.2
Farsi 1.6
French 1.6
Thai 1.2
Tagalog 0.9
German 0.7
Table 4: Share of all original tweets (dropping retweets) in TwitterDay by language. Retained languages are bolded.
Appendix B Annotation
B.1 Annotator demographics
We recruit a team of 30 annotators, that is three for each language and country. We provide detailed demographics for each language and country below:
Arabic
We recruit three Arabic-speaking male annotators. All three annotators are 18-29. Two annotators are educated to undergraduate level and the other one to taught masters. All three are native Arabic speakers respectively from Egypt, Tunisia and the United Arab Emirates.
English
We recruit two English-speaking female annotators and one male. Two annotators out of three are 18-29 while the other one is 30-39. One annotator is educated to undergraduate level, another one to taught masters and the last one to research degree (i.e. PhD). Two are native English speakers, respectively from the United States and Nigeria, and one is a non-native German citizen fluent in English.
French
We recruit two French-speaking male annotators and one female. One annotator out of three is 18-29 while the other two are 30-39. All annotators are educated to taught masters level. All three are native French speakers. Two annotators are French citizens while the last one is Canadian.
German
We recruit three German-speaking male annotators. Two annotators out of three are 18-29 while the other one is 30-39. Two annotators are educated to undergraduate level and the last one to research degree (i.e. PhD). All three are native German speakers, all from Germany.
Indonesian
We recruit two Indonesian-speaking male annotators and one female. Two annotators out of three are 18-29 while the other one is 30-39. Two annotators are educated to undergraduate level and the last one to research degree (i.e. PhD). All three are native Indonesian speakers, all from Indonesia.
Portuguese
We recruit two Portuguese-speaking female annotators and one male. All annotators are 30-39 and are educated to taught masters level. Two are native Portuguese speakers from Brazil and one is a non-native Portuguese-speaking Mexican citizen fluent in Portuguese from Portugal.
Spanish
We recruit two Spanish-speaking female annotators and one male. All annotators are 30-39. Two annotators are educated to undergraduate level and the last one to taught masters. All three are native Spanish speakers from Mexico.
Turkish
We recruit two Turkish-speaking male annotators and one female. Two annotators are 18-29 and the last one is 30-39. One annotator is educated to undergraduate level and the other two to taught masters. All three are native Turkish speakers from Turkey.
United States
We recruit two female annotators and one male. Two annotators out of three are 18-29 while the other one is 30-39. Two annotators are educated to taught masters and the last one to research degree (i.e. PhD). Two annotators are American citizens while the last one is a German citizen residing in the United States for more than five years.
India
We recruit two male annotators and one female. All three annotators are 18-29 and are educated to undergraduate level. All three annotators are Indian citizens.
Nigeria
We recruit two female annotators and one male. All three annotators are 30-39 and are educated to undergraduate level. All three annotators are Nigerian citizens.
Kenya
We recruit two female annotators and one male. All three annotators are 30-39 and are educated to undergraduate level. All three annotators are Kenyan citizens.
B.2 Annotation guidelines
B.3 Recruitment and compensation
Annotators were recruited through email outreach. Net hourly rate ranged from 5 to 24 USD depending on country of residence, education and experience.
Appendix C Academic Hate Speech Datasets
C.1 Language-level Supersets
The language-level supersets are documented in Tonneau et al. (2024a) and obtained from Hugging Face [^3] https://huggingface.co/collections/manueltonneau/hate-speech-supersets-664ef6d2bc40cce7a8b1092f. The number of annotated datapoints in these supersets for each language is:
  • Arabic: 464,260
  • English: 590,142
  • French: 18,071
  • German: 60,680
  • Indonesian: 14,306
  • Portuguese: 43,222
  • Spanish: 34,811
  • Turkish: 115,408
C.2 Country-level Dataset Survey
In the absence of supersets at the country-level, we conduct a survey of hate speech datasets grounded geographically in India, Nigeria and Kenya.
C.2.1 Surveying approach
To identify HS datasets, we rely on three data sources. First, we inspect the Hate Speech Data Catalogue [^4] https://hatespeechdata.com/ Vidgen and Derczynski (2020). Second, we inspect the datasets listed in the latest survey of HS datasets Poletto et al. (2021). Finally, we conduct a Google search for each country and inspect the links of the first three result pages in each case. We keep only the datasets that fit the following three criteria:
  1. 1.
    The dataset is documented, meaning it is attached to a research paper or a README file describing its construction.
  2. 2.
    The dataset is either publicly available or could be retrieved after contacting the authors.
  3. 3.
    The dataset focuses on HS, defined broadly as “any kind of communication in speech, writing or behavior, that attacks or uses pejorative or discriminatory language with reference to a person or a group on the basis of who they are, in other words, based on their religion, ethnicity, nationality, race, color, descent, gender or other identity factor” UN (2019).
C.2.2 Target categorization
We categorize each surveyed dataset in terms of the target categories they focus on. To do so, we look for mentions of target focus in the data documentation. If it is absent, we inspect the sampling approach and specifically the keywords used for sampling and use these to determine the target focus.
C.2.3 Surveyed datasets
We list below the retained datasets for each country, as well as their target focus:
India
  1. 1.
    A Dataset of Hindi-English Code-Mixed Social Media Text for Hate Speech Detection Bohra et al. (2018). Target focus: race/ethnicity/national origin, gender, caste, religion
  2. 2.
    Overview of the HASOC track at FIRE 2019: Hate Speech and Offensive Content Identification in Indo-European Languages Mandl et al. (2019). Target focus: religion, caste, gender, politics
  3. 3.
    Hostility Detection Dataset in Hindi Bhardwaj et al. (2020). Target focus: race/ethnicity/national origin, religion
  4. 4.
    Listening to Affected Communities to Define Extreme Speech: Dataset and Experiments Maronikolakis et al. (2022). Target focus: religion, caste
  5. 5.
    Uncovering Political Hate Speech During Indian Election Campaign: A New Low-Resource Dataset and Baselines Jafri et al. (2023). Target focus: politics
We identify two additional datasets Mathur et al. (2018); Saroj and Pal (2020) that we leave out of the analysis as we could not retrieve them.
Nigeria
  1. 1.
    HERDPhobia: A Dataset for Hate Speech against Fulani in Nigeria Aliyu et al. (2022). Target focus: race/ethnicity/national origin
  2. 2.
    Detection of Hate Speech Code Mix Involving English and Other Nigerian Languages Ndabula et al. (2023). Target focus: race/ethnicity/national origin, religion, politics
  3. 3.
    EkoHate: Abusive Language and Hate Speech Detection for Code-switched Political Discussions on Nigerian Twitter Ilevbare et al. (2024). Target focus: politics
  4. 4.
    NaijaHate: Evaluating Hate Speech Detection on Nigerian Twitter Using Representative Data Tonneau et al. (2024b). Target focus: race/ethnicity/national origin, gender, sexual orientation, religion
Kenya
  1. 1.
    Building and annotating a codeswitched hate speech corpora Ombui et al. (2021). Target focus: race/ethnicity/national origin, politics
  2. 2.
    Listening to Affected Communities to Define Extreme Speech: Dataset and Experiments Maronikolakis et al. (2022). Target focus: race/ethnicity/national origin, religion, politics
Appendix D Models
D.1 Hugging Face models
We list the Hugging Face models used as benchmark below for each language and country:
Arabic
  • IbrahimAmin-marbertv2-finetuned-
    egyptian-hate-speech-detection
    Ahmed et al. (2022)
  • Hate-speech-CNERG/dehatebert-mono-
    arabic
    Aluru et al. (2020)
English
  • Hate-speech-CNERG/bert-base-uncased-
    hatexplain
    Mathew et al. (2021)
  • facebook/roberta-hate-speech-
    dynabench-r4-target
    Vidgen et al. (2021)
  • Hate-speech-CNERG/dehatebert-mono-
    english
    Aluru et al. (2020)
  • IMSyPP/hate_speech_en Kralj Novak et al. (2022)
  • pysentimiento/bertweet-hate-speech Pérez et al. (2021)
French
  • Hate-speech-CNERG/dehatebert-mono-
    french
    Aluru et al. (2020)
  • Poulpidot/distilcamenbert-french-
    hate-speech
  • julio2027/French_hate_speech_
    CamemBERT_v3
German
  • jagoldz/gahd Goldzycher et al. (2024)
  • deepset/bert-base-german-
    cased-hatespeech-GermEval18Coarse
  • jorgeortizv/BERT-hateSpeech
    Recognition-German
  • Hate-speech-CNERG/dehatebert-
    mono-german
    Aluru et al. (2020)
  • shahrukhx01/gbert-hasoc-german-2019
Indonesian
  • Hate-speech-CNERG/dehatebert-mono-
    indonesian
    Aluru et al. (2020)
Portuguese
  • knowhate/HateBERTimbau-yt-tt
  • Hate-speech-CNERG/dehatebert-mono-
    portugese
    Aluru et al. (2020)
Spanish
  • Hate-speech-CNERG/dehatebert-mono
    -spanish
    Aluru et al. (2020)
  • pysentimiento/robertuito-hate-speech Pérez et al. (2021)
  • jorgeortizfuentes/chilean-spanish-
    hate-speech
Turkish
  • ctoraman/hate-speech-berturk Toraman et al. (2022)
United States
We use the same models as for English for the United States.
India
  • Hate-speech-CNERG/kannada-codemixed-
    abusive-MuRIL
    Das et al. (2022)
  • Hate-speech-CNERG/marathi-codemixed-
    abusive-MuRIL
    Das et al. (2022)
  • Hate-speech-CNERG/bengali-abusive-
    MuRIL
    Das et al. (2022)
  • Hate-speech-CNERG/tamil-codemixed-
    abusive-MuRIL
    Das et al. (2022)
  • Hate-speech-CNERG/english-abusive-
    MuRIL
    Das et al. (2022)
Nigeria
  • worldbank/naija-xlm-twitter-base-hate Tonneau et al. (2024b)
Kenya
We have not found any hate speech detection models tailored to the Kenyan context on Hugging Face.
D.2 Zero-shot learning
The prompt we use for zero-shot learning: “Here is a social media post: [POST]. Classify this post as “hateful” or “non-hateful”. Only return the predicted class.”.
The prompt is run three times for each tweet. We then define the hateful score as the share of the three times for which the model predicted that the tweet was hateful. We then use this score to compute the average precision. We use all default values for the main hyperparameters, including 1 for temperature.
D.3 Computing infrastructure
For inference, we used either V100 (32GB) or RTX8000 (48GB) GPUs.
D.4 Number of parameters
The supervised learning models are largely based on BERT base which has 110 million parameters. The decoder-based models (Aya and Llama3.1) both have 8 billion parameters.
Appendix E Additional Results
E.1 Figures
We provide additional figures representing respectively the precision-recall curves across languages and countries (Figure 5), the comparison between target-level academic focus and target-level share of all hate in HateDay (Figure 6) and the relationship between the offensive-to-hateful count ratio and the share of offensive content in top-scored tweets of HateDay (Figure 7).
E.2 Individual open-source model performance
We provide detailed average precision results on the three evaluation sets for each individual open-source model (see D.1 for the full list) and across languages (Table 5) and countries (Table 6).
Language Model AD HC HD
Arabic IbrahimAmin/marbertv2-finetuned-
egyptian-hate-speech-detection 25.8±0.6 76.9 $\pm$2.0 5.8$\pm$3.0
Hate-speech-CNERG-
dehatebert-mono-arabic
16.4±0.4 83.1$\pm$1.7 1.6$\pm$1.4
Perspective API 18.7±0.4 89.6$\pm$1.1 10.2$\pm$5.6
English Hate-speech-CNERG/bert-base-uncased-hatexplain 47.2±0.3 74.3±2.0 2.5±1.0
facebook/roberta-hate-speech-dynabench-r4-target 55.3±0.3 98.7±0.4 4.2±2.2
Hate-speech-CNERG/dehatebert-mono-english 48.7±0.3 75.0±2.0 7.2±4.4
IMSyPP/hate_speech_en 21.7±0.1 65.9±2.0 0.6±0.2
pysentimiento/bertweet-hate-speech 38.2±0.3 82.2±1.7 9.0±3.9
Perspective API 52.9±0.3 95.1±0.6 10.1±3.5
French Hate-speech-CNERG/dehatebert-mono-french 27.5±1.0 74.5±1.9 2.7±0.6
Poulpidot/distilcamenbert-french-hate-speech 27.6±0.9 87.2±1.5 4.2±0.9
julio2027/French_hate_speech_CamemBERT_v3 35.7±1.3 89.9±1.2 16.0±4.2
Perspective API 40.1±1.4 96.9±0.4 37.7±6.0
German jagoldz-gahd 64.3±1.2 97.2±0.5 19.3±5.0
deepset/bert-base-german-cased-hatespeech-
GermEval18Coarse 22.6±0.8 82.1±1.9 8.5±2.6
jorgeortizv/BERT-hateSpeechRecognition-German 19.9±0.7 76.8±2.0 3.4±1.1
Hate-speech-CNERG/dehatebert-mono-german 18.1±0.7 80.2±1.9 1.8±0.5
shahrukhx01-gbert-hasoc-german-2019 19.0±0.7 79.4±1.8 5.6±1.4
Perspective API 53.6±1.1 96.0±0.5 18.8±4.4
Indonesian Hate-speech-CNERG-dehatebert-mono-indonesian 90.2±0.7 - 3.5±2.3
Perspective API 65.5±1.3 - 11.1±5.9
Portuguese knowhate/HateBERTimbau-yt-tt 18.7±0.9 85.8±1.6 3.1±1.7
Hate-speech-CNERG/dehatebert-mono-portugese 19.1±1.0 74.4±1.8 1.3±0.5
Perspective API 30.9±1.5 94.6±0.7 14.5±6.9
Spanish Hate-speech-CNERG/dehatebert-mono-spanish 60.3±1.0 78.2±1.9 3.3±1.0
pysentimiento/robertuito-hate-speech 72.6±0.9 87.0±1.5 8.2±1.6
jorgeortizfuentes/chilean-spanish-hate-speech 54.9±1.1 82.3±1.6 11.6±3.3
Perspective API 50.9±1.1 96.0±0.5 34.1±6.2
Turkish ctoraman/hate-speech-berturk 32.1±0.4 - 1.9±0.6
Perspective API 31.4±0.4 - 6.5±3.8
Table 5: Detailed open-source and Perspective model performance across languages and evaluation sets, as measured by average precision (%). Metrics are reported with 95% bootstrapped confidence intervals. We report performance on three evaluation sets: academic hate speech datasets (AD) combined for a given language, HateCheck functional tests (HC) and HateDay (HD). HC does not cover Indonesian and Turkish.
Country Model AD HD
United States Hate-speech-CNERG/bert-base-uncased-hatexplain 47.2±0.3 4.6±2.3
facebook/roberta-hate-speech-dynabench-r4-target 55.3±0.3 3.1±0.9
Hate-speech-CNERG/dehatebert-mono-english 48.7±0.3 4.6±2.3
IMSyPP/hate_speech_en 21.7±0.1 0.6±0.2
pysentimiento/bertweet-hate-speech 38.2±0.3 8.2±2.7
Perspective API 52.9±0.3 12.3±3.4
India Hate-speech-CNERG/kannada-codemixed-abusive-MuRIL 48.6±1.1 5.9±1.1
Hate-speech-CNERG/marathi-codemixed-abusive-MuRIL 43.4±1.0 3.7±0.9
Hate-speech-CNERG/bengali-abusive-MuRIL 40.0±1.0 3.3±0.7
Hate-speech-CNERG/tamil-codemixed-abusive-MuRIL 47.1±1.0 6.2±1.2
Hate-speech-CNERG/english-abusive-MuRIL 54.3±1.1 7.8±1.6
Perspective API 61.9±1.1 42.9±5.8
Nigeria worldbank/naija-xlm-twitter-base-hate 65.7±1.4 30.9±1.3
Perspective API 44.1±1.6 8.6±6.5
Kenya Perspective API 31.6±0.9 9.1±6.1
Table 6: Detailed open-source and Perspective model performance across countries and evaluation sets, as measured by average precision (%). Metrics are reported with 95% bootstrapped confidence intervals. We report performance on two evaluation sets: academic hate speech datasets (AD) combined for a given language and HateDay (HD).
E.3 Human-in-the-loop moderation
In the human-in-the-loop setting, we compute the proportion of all tweets to be reviewed by human annotators as follows. We first compute the number of predicted positives PP:
$PP=TP+FP=TP/\textrm{precision}$
$PP=(\textrm{recall}*(TP+FN))/\textrm{precision}$
$PP=(\textrm{recall}*\textrm{total \# hateful tweets})/\textrm{precision}$
$PP=(\textrm{recall}*\textrm{base rate}*\textrm{total \# tweets})/\textrm{precision}$
We then derive the share $S$ of all tweets that are predicted positive by a given model, that is the share of all tweets that will be reviewed by human moderators in a human-in-the-loop approach, by dividing $PP$ by the total number of tweets:
$S=(\textrm{recall}*\textrm{base rate})/\textrm{precision}$
with the base rate equal to the prevalence of hateful content in HateDay for that specific language. We finally use the precision and recall values from the precision-recall curve to derive the curve illustrating the relationship between recall and $S$.
Figure 5: Precision-recall curves for each language and country
Figure 6: Comparison between the target-level share of all hate in HateDay and the share of all hate speech datasets for each language or country and target combination. The language-level data for hate speech datasets is taken from Yu et al. (2024).
Figure 7: Offensive-to-hateful count ratio versus the share of offensive content in top tweets (in %). Top tweets are defined as the 1% (N=200) top-scored tweets in HateDay using the best performing model on the latter.

### v2

@@
 Abstract
-To tackle the global challenge of online hate speech, a large body of research has developed detection models to flag hate speech in the sea of online content.
-Yet, due to systematic biases in evaluation datasets, detection performance in real-world settings remains unclear, let alone across geographies.
-To address this issue, we introduce HateDay, the first global hate speech dataset representative of social media settings, randomly sampled from all tweets posted on September 21, 2022 for eight languages and four English-speaking countries.
-Using HateDay, we show how the prevalence and composition of hate speech varies across languages and countries.
-We also find that evaluation on academic hate speech datasets overestimates real-world detection performance, which we find is very low, especially for non-European languages.
-We identify several factors explaining poor performance, including models’ inability to distinguish between hate and offensive speech, and the misalignment between academic target focus and real-world target prevalence.
-We finally argue that such low performance renders hate speech moderation with public detection models unfeasible, even in a human-in-the-loop setting which we find is prohibitively costly.
-Overall, we emphasize the need to evaluate future detection models from academia and platforms in real-world settings to address this global challenge.
+To address the global challenge of online hate speech, prior research has developed detection models to flag such content on social media.
+However, due to systematic biases in evaluation datasets, the real-world effectiveness of these models remains unclear, particularly across geographies.
+We introduce HateDay, the first global hate speech dataset representative of social media settings, constructed from a random sample of all tweets posted on September 21, 2022 and covering eight languages and four English-speaking countries.
+Using HateDay, we uncover substantial variation in the prevalence and composition of hate speech across languages and regions.
+We show that evaluations on academic datasets greatly overestimate real-world detection performance, which we find is very low, especially for non-European languages.
+Our analysis identifies key drivers of this gap, including models’ difficulty to distinguish hate from offensive speech and a mismatch between the target groups emphasized in academic datasets and those most frequently targeted in real-world settings.
+We argue that poor model performance makes public models ill-suited for automatic hate speech moderation and find that high moderation rates are only achievable with substantial human oversight.
+Our results underscore the need to evaluate detection systems on data that reflects the complexity and diversity of real-world social media.
 Content warning: This article contains illustrative examples of hateful content.
 1 Introduction
-Social media users often encounter hate speech in their feeds UNESCO (2023), fueling fears of its impacts on offline violence Müller and Schwarz (2021).
-To counter this phenomenon, a mature body of research has developed automatic hate speech detection models Vidgen and Derczynski (2020).
-Yet, past work in this space suffers from three major shortcomings.
-Figure 1: HateDay consists of twelve annotated representative sets (N=20K each) randomly sampled from all tweets posted on September 21, 2022 in eight languages (Arabic, English, French, German, Indonesian, Portuguese, Spanish, and Turkish) and originating from four countries where English is the main language on Twitter (United States, India, Nigeria, Kenya).
-First, the performance of hate speech detection systems in real-world settings remains largely unclear.
-On the one hand, detection models used in industry are not public and detection performance metrics are usually missing from platform transparency reports or sometimes misleading Giansiracusa (2021).
-On the other hand, evaluation of hate speech models in academic research has been done on biased datasets whose characteristics greatly differ from real-world settings, from class imbalance to topic distribution Nejadgholi and Kiritchenko (2020), raising concerns about an overestimation of the reported detection performance Arango et al.
+Social media users frequently encounter hate speech in their feeds UNESCO (2023), raising concerns about its potential to incite offline violence Müller and Schwarz (2021).
+In response, a substantial body of research has focused on developing automated hate speech detection models Vidgen and Derczynski (2020).
+However, prior work in this area suffers from three major limitations.
+Figure 1: HateDay consists of twelve annotated representative sets (N=20K each) randomly sampled from all tweets posted on September 21, 2022.
+The dataset covers eight languages (Arabic, English, French, German, Indonesian, Portuguese, Spanish, and Turkish) and four countries where English is the main language on Twitter (India, Kenya, Nigeria, and the United States).
+First, the performance of hate speech detection systems in real-world social media settings remains largely unknown.
+Industry models are typically not publicly available, and platform transparency reports omit detailed performance metrics or present them in ways that can be misleading or difficult to interpret Giansiracusa (2021).
+Meanwhile, academic evaluations rely on biased datasets that diverge considerably from real-world distributions, especially regarding class imbalance and topic diversity Nejadgholi and Kiritchenko (2020).
+This raises concerns that reported performance may be substantially overestimated and not generalize well outside such controlled settings Arango et al.
 (2019); Wiegand et al.
 (2019).
-The second major limitation of academic research on automatic detection of hate speech lies in its limited linguistic scope, with most resources developed for English Poletto et al.
+The second key limitation of academic research on automatic hate speech detection lies in its narrow language focus.
+Indeed, most available datasets and detection models have been developed for English Poletto et al.
 (2021); Tonneau et al.
 (2024a).
-While such unequal resource allocation is likely to result in unequal performance across languages, the use of customized datasets for evaluation renders cross-lingual performance comparisons unfeasible.
-Finally, the focus on languages in the development of hate speech detection systems may hide performance differences between countries that share a common language Ghosh et al.
+While such unequal resource allocation is likely to result in unequal performance across languages, the widespread use of customized datasets for evaluation hinders meaningful cross-lingual comparisons.
+Finally, focusing on languages in hate speech detection can hide performance differences between countries sharing a language Ghosh et al.
 (2021).
-While recent work has focused on developing datasets grounded in geographical location rather than language Maronikolakis et al.
+Although recent work develops datasets based on geography rather than language Maronikolakis et al.
 (2022); Castillo-lópez et al.
-(2023), cross-country performance differences remain largely unclear.
-In this context, it is necessary to develop multilingual evaluation sets that are representative of real-world settings and that provide a comparable setting across languages and countries with a common language.
-In this article, we aim to address these three shortcomings and provide an evaluation of state-of-the-art hate speech detection performance on real-world social media data across languages and countries.
-For this purpose, we introduce HateDay, a global Twitter dataset consisting of 240,000 tweets sampled from all tweets posted on September 21, 2022 and annotated for hate speech detection.
-Specifically, we draw random samples of 20,000 tweets each for eight popular languages on Twitter: Arabic, English, French, German, Indonesian, Portuguese, Spanish, and Turkish.
-In addition, we draw random samples of 20,000 tweets each for four countries whose major language on Twitter is English: the United States, India, Nigeria, and Kenya.
-We show that the prevalence and composition of hate varies across languages and countries.
-We then use HateDay to evaluate public hate speech detection models across languages and countries.
-We find that hate speech detection performance is largely overestimated by evaluation on academic hate speech data from prior work and is very low in real-world social media settings, particularly for non-European languages.
-We identify several factors explaining such poor model performance including the inability of models to distinguish between hate and offensive speech, and the misalignment between academic target coverage and real-world target prevalence.
-Finally, in light of our performance results, we estimate the feasibility of hate speech moderation using publicly available models.
-We argue that automatic moderation alone is too error-prone to be desirable and show that human-in-the-loop moderation, whereby models flag potentially hateful content for human verification, proves prohibitively costly across languages and countries, despite its potential for improved accuracy.
+(2023), cross-country performance variations remain unclear.
+Addressing this limitation requires evaluation datasets providing a comparable evaluation setting across countries with a common language.
+In this article, we address these three limitations by evaluating the performance of state-of-the-art hate speech detection models on real-world social media data across multiple languages and countries.
+We introduce HateDay, a global Twitter dataset which consists of 240,000 tweets, randomly sampled from all tweets posted on September 21, 2022 and annotated for hate speech (Figure 1, left).
+Specifically, we sample 20,000 tweets each for eight widely used languages on Twitter: Arabic, English, French, German, Indonesian, Portuguese, Spanish, and Turkish.
+We also sample 20,000 tweets each for four English-speaking countries with a sizable Twitter presence: India, Kenya, Nigeria, and the United States.
+Using HateDay, we provide insights on real-world online hate (Figure 1, right).
+We show that the prevalence and composition of hate vary significantly across languages and countries.
+We then evaluate publicly available hate speech detection models across this global landscape.
+Our findings reveal that detection performance is substantially overestimated when assessed on standard academic datasets and is strikingly low in real-world social media settings—especially for non-European languages.
+We identify several factors contributing to this poor performance, including models’ limited ability to distinguish hate from offensive speech, and a mismatch between the target categories prevalent in academic datasets and those prevalent in real-world discourse.
+Given these results, we examine the feasibility of using public models for hate speech moderation.
+We argue that automatic moderation—where posts flagged as hateful are moderated without human oversight—is currently too error-prone to be a viable approach.
+In contrast, we find that human-in-the-loop systems, in which flagged content is reviewed by moderators, can achieve higher moderation rates across languages and countries, albeit at the cost of substantial human review.
 In sum, our contributions are:
   1. 1.
-    HateDay, the first global hate speech dataset representative of social media settings [^1] Available at https://huggingface.co/datasets/manueltonneau/hateday, composed of 240,000 annotated tweets covering eight languages and four countries.
+    HateDay, the first global hate speech dataset representative of social media settings, composed of 240,000 annotated tweets covering eight languages and four countries [^1] Available at https://huggingface.co/datasets/manueltonneau/hateday
   2. 2.
-    a comparison of hate speech prevalence and composition across languages and countries
+    a cross-lingual and cross-national comparison of hate prevalence and composition
   3. 3.
-    an estimation of cross-lingual and geographic hate speech real-world detection performance, including a detailed analysis of pitfalls of current detection models
+    an evaluation of real-world detection performance across languages and countries, along with a detailed analysis of the limitations of current detection models
   4. 4.
-    an estimation of the feasibility of hate speech moderation using publicly available models
+    an assessment of the feasibility of hate speech moderation using publicly available models, including a comparison of automatic and human-in-the-loop approaches
+2 The HateDay Dataset
 Figure 2: Prevalence of harmful content (left) and targets of hate speech (right) in HateDay.
 The target category “Other” contains rare target labels such as “Caste”, “Age”, “Occupation”, “Disability” and “Social Class”.
-2 The HateDay Dataset
 2.1 Data Collection
@@
 This amounts to approximately 375 million tweets posted by 40 million users platform-wide.
-We filter TwitterDay both at the language and country levels to enable a more structured analysis.
+We filter TwitterDay both at the language and country levels.
 At the language level, we focus on the eight most popular languages in the TwitterDay dataset for which there exist hate speech detection resources (details in §A.1), namely Arabic, English, French, German, Indonesian, Portuguese, Spanish and Turkish.
-To assess differences between countries with a common language, we also filter at the country-level, focusing on four countries for which English is the main language on Twitter, namely the United States, India, Nigeria and Kenya.
-For locating users in each country, we use the Google Geocoding API to infer user country location based on their self-provided profile location Hecht et al.
+To assess differences between countries with a common language, we also filter at the country-level, focusing on four countries for which English is the main language on Twitter, namely India, Kenya, Nigeria, and the United States.
+We use the Google Geocoding API to infer user country location based on their self-provided profile location, knowing this has limitations Hecht et al.
 (2011).
 We drop retweets to focus on original content and randomly sample 20,000 tweets for each of the eight retained languages and each of the four countries, ensuring that each language-specific and country-specific sample is representative of real-world Twitter settings in that particular language or country.
-The HateDay dataset corresponds to the combination of all the annotated random samples for each language and country, for a total of 240,000 annotated tweets.
+The HateDay dataset corresponds to the combination of all annotated random samples for each language and country, totaling 240,000 annotated tweets.
 2.2 Data Annotation
 We recruit a team of 36 annotators, three per language or country.
-For languages that are spoken widely around the globe, such as English, Arabic or French, we maximize the diversity of annotator origins where possible (details in §B.1).
-We follow a prescriptive approach to manage annotator subjectivity in our annotation task Rottger et al.
-(2022) by instructing annotators to strictly adhere to extensive annotation guidelines that describe our taxonomy of hate speech (provided in §B.2).
+For languages that are spoken widely around the globe, such as English, Arabic or French, we maximize the diversity of annotator origins as much as possible (details in §B.1).
+We follow a prescriptive approach to manage annotator subjectivity in our annotation task Röttger et al.
+(2022b) by instructing annotators to strictly adhere to extensive annotation guidelines that describe our taxonomy of hate speech (provided in §B.2).
 Following prior work Davidson et al.
 (2017); Mathew et al.
-(2021), the annotation consists of assigning each tweet to one of three classes: (i) hateful, if it promotes violence or contains a direct attack, abuse or threat against an individual or a group based on the perceived belonging of a certain characteristic (e.g., gender, race), (ii) offensive, if it contains any form of non-acceptable language that is not hateful, including insults, threats, and posts containing profane language or swear words Zampieri et al.
+(2021), the annotation consists of assigning each tweet to one of three classes: (i) hateful, if it promotes violence or contains a direct attack, abuse or threat against an individual or a group based on the perceived belonging of a certain characteristic (e.g., gender, race), (ii) offensive, if it contains potentially objectionable language that is not hateful, including insults, threats, and posts containing profanity or swear words Zampieri et al.
 (2019), or (iii) neutral if it is neither hateful nor offensive.
-For tweets labeled as hateful, we also ask annotators to specify the groups being targeted.
+For tweets labeled as hateful, we also ask annotators to specify the groups being targeted as a free-text label.
 We provide examples of tweets for each class in Table 1.
-Class Target Category Examples Origin
+Class Target Category Examples Split
 Politics Greens are vermin German
@@
 Table 1: Examples of tweets for each class and main target categories. Offensive tweets, by definition, have no target. English translations are displayed in italic.
-For each language and country, we conduct a pilot annotation to train annotators on the task.
-Specifically, we have each annotator label 100 posts sampled from a collection of all annotated hate speech datasets Tonneau et al.
-(2024a) open-sourced on Hugging Face, a major open source platform.
-We repeat this task until the inter-annotator agreement, measured by Krippendorff’s $\alpha$, reaches 0.7.
-After the pilot, we task each annotator to label the random sample of 20,000 tweets in their respective language or country.
-Each tweet is labeled by three annotators, and the final label is determined by majority vote.
-To minimize the amount of hate that is wrongly categorized as not hate, all tweets without a majority vote or with at least one hateful label out of three are re-examined by all three annotators and a task expert to derive their final label.
-Across all languages and countries, the three annotators agreed on 93.4% of all labeled tweets, two out of three agreed in 6.4% of cases, and all three of them disagreed in 0.2% of cases.
+For each language and country, we conduct a pilot annotation phase to train annotators on the task.
+Specifically, we have each annotator label 100 posts, sampled from a collection of all annotated hate speech datasets Tonneau et al.
+(2024a), which have been open-sourced on Hugging Face.
+We repeat this task until the inter-annotator agreement, measured by Krippendorff’s $\alpha$, reaches a threshold of 0.7.
+After completing the pilot, we then task each annotator with labeling the random sample of 20,000 tweets in their respective language or country.
+Each tweet is labeled by three annotators, and the final label is determined by a simple majority vote.
+Across all languages and countries, the three annotators agreed on 93.4% of all labeled tweets; two out of three agreed in a further 6.4% of cases, and all three disagreed in only 0.2% of cases.
 2.3 Descriptive Statistics
 Prevalence of harmful content
-We find that the prevalence of hate speech on Twitter on the day of analysis is very low, representing less than 2% of all posts for all considered languages and countries and with an average prevalence of 0.7% (Figure 2, left).
-In contrast, we observe that offensive content is much more prevalent than hate speech across all languages and countries, ranging from 2.5 times more prevalent for French to 19 times more prevalent for Portuguese.
-We also observe large differences in the prevalence of hate speech between languages and countries.
+We find that the prevalence of hate speech on Twitter on the day of analysis is very low, representing less than 2% of all posts across languages and countries, with an average prevalence of 0.7% (Figure 2, left).
+In contrast, offensive content is substantially more prevalent than hate speech across all considered languages and countries, ranging from 2.5 times more prevalent for French to 19 times more prevalent for Portuguese.
+We also find notable differences in the prevalence of hate speech between languages and countries.
 For instance, 1.0% of German tweets are hateful whereas only 0.3% of Indonesian tweets are hateful.
-We observe similar gaps at the country level, with the share of hateful posts being much higher in India (1.7%) than in Kenya (0.2%) or Nigeria (0.3%).
+We observe similar gaps at the country level, with the share of hateful posts being much higher in India—1.7%—compared to 0.2% in Kenya and 0.3% in Nigeria.
 Main targets of hate speech
@@
 We also find notable differences across languages and countries.
-For instance, political hate speech is prevalent in English, German, Indonesian, Portuguese, and Spanish, as well as in the U.S.
-national context, whereas it is less present in Turkish and Nigerian tweets, and not present at all in Arabic tweets.
-Also, religious hate speech, mostly in the form of Islamophobia, dominates in India whereas it is less prominent elsewhere and completely absent from the sample in the Kenyan case.
+For instance, political hate speech is prevalent in English, German, Indonesian, Portuguese, and Spanish, as well as in the US national context, representing up to 66% of all hate in the German context.
+In contrast, it is less present in Turkish and Nigerian tweets, and not present at all in Arabic tweets.
+Also, religious hate speech represents 41% of all hate in India, mostly in the form of Islamophobia, whereas it is less prominent elsewhere and completely absent from the Kenyan sample.
 Finally, we note that some forms of hate speech are unique to specific contexts, such as casteism in India.
@@
 (2022), a widely used toxic language detection system also based on supervised learning.
+Specifically, we use the API’s “Identity Attack” attribute, as it most closely aligns with our definition of hate speech.
 Zero-shot learning
@@
 (2024) for zero-shot learning.
-We do so because Aya is designed to be multilingual, and Llama3.1 is one of the most capable open-source decoder-only models available at the time of the analysis.
-We use smaller architectures due to computing constraints and provide the prompt for zero-shot learning in §D.2.
+We do so because Aya is designed to be multilingual, and Llama3.1 is one of the most capable open-source models available at the time of our analysis.
+We use small model sizes due to compute constraints.
+For the zero-shot prompt, see §D.2.
 3.2 Evaluation
 We evaluate models on the HateDay (HD) datasets to estimate real-world detection performance.
-Additionally, we compare the latter with the performance on two traditional evaluation methods, namely academic hate speech datasets and functional tests.
+For comparison, we also test models in two traditional evaluation settings, namely on academic hate speech datasets and functional tests.
 Academic hate speech datasets (AD)
-We measure performance on academic hate speech datasets to understand how results from past work generalize to our more realistic HateDay datasets.
-We rely on open-sourced hate speech supersets combining all existing hate speech datasets for all eight languages of interest Tonneau et al.
+We measure performance on academic hate speech datasets to understand how results from past work generalize to more realistic settings.
+We rely on supersets combining all existing hate speech datasets for all eight languages of interest Tonneau et al.
 (2024a).
 Given that English hate speech datasets mostly originate from the US Tonneau et al.
-(2024a), we use English hate speech datasets both for evaluating English-centered and US-centered models.
+(2024a), we use English hate speech datasets both for evaluating English- and US-centered models.
 In the absence of supersets for India, Nigeria and Kenya, we survey all existing datasets for each country and combine them to build the supersets (details in §C.2).
-We restrict the evaluation to a 10% random sample of all academic hate speech datasets for each language and country to limit the computing resources required for inference (details in §C.1).
+We restrict the evaluation to a 10% random sample of all academic datasets for each language and country to limit inference costs (details in §C.1).
 Functional tests (HC)
-We measure performance on functional tests to estimate models’ ability to handle known challenges in hate speech detection.
+We measure performance on functional tests to estimate the ability of models to handle known challenges in hate speech detection.
 We use HateCheck Röttger et al.
-(2021, 2022), a suite of functional tests for hate speech detection models, which covers six of eight languages in HateDay, missing Indonesian and Turkish.
-We evaluate model performance on each dataset using average precision, which corresponds to the area under the precision-recall curve, and is well suited when class imbalance is high, as is the case in HateDay.
+(2021, 2022a), a suite of functional tests for hate speech detection models, which covers six of eight languages in HateDay, missing Indonesian and Turkish.
+Evaluation metric
+We evaluate model performance using average precision, which corresponds to the area under the precision-recall curve, and is well suited when class imbalance is high.
 4 Results
@@
 Best OS 25.8 76.9 5.8 38.2 82.2 9 35.7 89.9 16 64.3 97.2 19.3 90.2 - 3.5 18.7 85.8 3.1 54.9 82.3 11.6 32.1 - 1.9
-Perspective 18.7 89.6 10.2 58.9 95.1 10.1 40.1 96.9 37.7 53.6 96 18.8 65.5 - 11.1 30.9 94.6 14.5 50.9 96 34.1 $31.4^{*}$ $\textbf{6.5}^{*}$
+Perspective 18.7 89.6 10.2 58.9 95.1 10.1 40.1 96.9 37.7 53.6 96 18.8 65.5 - 11.1 30.9 94.6 14.5 50.9 96 34.1 31.4* - 6.5*
 Table 2: Model performance across languages and evaluation sets, as measured by average precision (%).
@@
 Performance across datasets
-Our most striking finding is that across the wide range of models, languages and countries considered in this study, the average precision on our representative HateDay data (HD) is just 9.4% on average, compared to 40% on the sample of public hate speech datasets (AD) and 87.2% on the functional tests (HC).
+Our most striking finding is that across the wide range of models, languages, and countries considered, detection performance is much lower on our representative HateDay dataset (HD) than on academic hate speech datasets (AD) or functional tests (HC).
+Indeed, average precision is just 9.4% on HD, compared to 40% on AD and 87.2% on HC.
 Performance across model types
 We find that supervised learning consistently outperforms zero-shot learning across almost all combinations of language, country, and evaluation sets.
-Indeed, the best open-source model on HateDay has an average precision of 36.0% on average across all combinations, whereas Llama3.1 8B, which performs on par with Aya23 8B, scores 28.4% on average.
-Additionally, despite being originally developed to detect toxic rather than hateful content, we observe that the Perspective API (39.1% average precision on average) outperforms open-source hate speech detection models in all languages but German.
-Still, the dominance of Perspective is more limited at the country-level where it performs best in the US and Indian contexts but not in the Nigerian context, and only in the Kenyan context because there are no supervised models to compare to.
+Indeed, the best open-source model on HateDay has an average precision of 41.1% on average across all combinations of language or country and dataset, whereas Aya23 8B, which performs on par with Llama3.1 8B, scores 32.4% on average.
+Additionally, despite being originally developed to detect toxic rather than hateful content, we observe that the Perspective API (44.3% average precision on average) outperforms open-source hate speech detection models in all languages except German.
+At the country-level, open-source models perform best in Nigeria, but Perspective API has higher performance in the US and India.
+We did not find open-source supervised models to compare Perpective API to in the Kenyan context.
 Model Type United States India Nigeria Kenya
@@
 Perspective 58.9 12.3 61.9 42.9 44.1 8.6 31.6 9.1
-Table 3: Model performance across countries and evaluation sets, as measured by average precision (%). There is no open-source model specifically for Kenya.
+Table 3: Model performance across countries and evaluation sets, as measured by average precision (%). There are no open-source models specifically for Kenya.
 Cross-lingual and geographic gaps
 We observe substantial performance differences in the real-world setting of HateDay across languages and countries.
-At the language-level, we find that the best average precision for European languages (23.1% on average), such as French, Spanish or German, is substantially higher than for non-European languages (9.3%), such as Arabic, Indonesian and Turkish.
-At the country-level, the average precision is by far highest on Indian tweets (42.9%) followed by Nigeria (30.9%), whereas the performance is rather low in both the U.S.
-(12.3%).
-and Kenyan (9.1%) contexts.
-4.2 Performance Analysis
-In the following, we conduct a more in-depth analysis into potential explanations for the poor hate speech detection performance as well as the cross-geographic performance gaps that we observe in real-world settings in §4.1.
+At the language level, average precision is higher for European languages—English, French, Spanish, Portuguese, and German—averaging 23.1%—than for non-European languages—Arabic, Indonesian, and Turkish—at 9.3% on average.
+At the country level, performance is highest on Indian tweets (42.9%), followed by Nigeria (30.9%), and lower in both the US (12.3%) and Kenya (9.1%).
+4.2 Reasons for Low Performance
+In the following, we conduct a more in-depth analysis into potential explanations for the poor hate speech detection performance as well as the cross-geographic performance gaps we observed in real-world settings in §4.1.
 Low precision and recall
 As the average precision corresponds to the area under the precision-recall curve, the same average precision may correspond to very different patterns in precision and recall.
-We therefore inspect the precision-recall curves of the best-performing models on HateDay for each language and country.
-We find that, while precision or recall may be higher depending on the context, both values remain very low and there are no situations where both values are above 50%, with the highest F1-score at 47% for India (Figure 5 in the Appendix).
+We therefore inspect the precision-recall curves of the best-performing models on HateDay for each language and country (Figure 5 in the Appendix).
+We find that, while precision or recall may be higher depending on the context, both values remain very low and there are no situations where both values are above 50%, with the highest F1-score at 47% for India.
 This highlights the role of both false positives and negatives in the observed low performance.
@@
 Figure 3: Share of hateful, offensive and neutral content in the top 5% scored tweets (N=1,000) in HateDay for each language and country.
-We use the score distribution of the best performing model on HateDay (Tables 2 and 3).
+We use the hatefulness score distribution of the best performing model on HateDay (Tables 2 and 3).
 The x-axis corresponds to the descending rank and each bar shows the distribution of content in a window of 50 tweets.
 Offensive false positives
-We find that offensive content constitutes a substantial share of the top of the score distribution — 32.3% in the top 50 (0.25%) scored tweets on average — for each language and country (Figure 3), thereby hurting precision.
-For Arabic, English, Portuguese and the U.S.
-and Kenyan national contexts, offensive content even appears more frequently than hate speech at the top of the score distribution.
-This problem is further aggravated by the fact that offensive content is more prevalent than hate speech on the day of the analysis (Figure 2).
-Indeed, we observe that as the prevalence ratio of offensive content to hate speech increases, the share of offensive content also gradually increases (Figure 7 in the Appendix).
+We find that offensive content constitutes a substantial share of the top of the hatefulness score distribution—32.3% of the top 50 (0.25%) scored tweets on average—for each language and country (Figure 3), thereby significantly hurting precision.
+For Arabic, English, Portuguese, and the US and Kenyan national contexts, offensive content even appears more frequently than hate speech at the very top of the score distribution.
+This problem is further aggravated by the fact that offensive content is more prevalent than hate speech (Figure 2).
+Indeed, we observe that as the prevalence ratio of offensive content to hate speech steadily increases, the share of offensive content at the top of the score distribution also gradually increases (Figure 7 in the Appendix), exacerbating the negative impact on precision.
 Qualitative analysis of false positives
-We further examine top-scored neutral tweets to identify other prominent features of false positives.
-We find that many such tweets contain mentions of hate speech (e.g., “wasn’t it the political scientist saying that Ukrainians are “subhumans”?”; “for them, women are like joints, they smoke them until they get high”).
-We also find several cases of statements that are not hateful but that would become hateful with a few changes of letters (e.g., “emigration is a catastrophy”; “this was noggers, man [^2] Portuguese original is: “isso foi noggers mano.”.
-“Noggers” means “disappointing” in Portuguese gamer slang.).
-Finally, we find that a substantial share of such tweets may be hateful but lack the context to fully determine their hatefulness.
-Such tweets often are replies and contain target community names (e.g., “@USER No, illegal migrants, they say it themselves”; “@USER We know you are gay”).
+We further examine top-scored neutral tweets to identify other prominent features of false positives beyond offensiveness.
+We find that many such tweets contain mentions of hate speech (e.g., “was it the political scientist saying that Ukrainians are ‘subhumans’?”).
+We also find several cases of statements that are not hateful but would be with a few changes of letters (e.g., “emigration is a catastrophy”).
+Finally, we find that a substantial share of such tweets, usually replies, may be hateful but lack the context to conclude (e.g., “@USER illegal migrants, they say it themselves”).
 Analysis of false negatives
-Next, we analyze false negatives, defined as hateful examples that do not appear in the top 1% of the best model’s score distribution for each language and country.
-We first examine the targets of these tweets to determine how their representation in false negatives compares to their overall prevalence in hate speech for each specific language or country.
-We find that politics and gender are overrepresented in false negatives, respectively in 10/12 and 7/12 of the studied language and country-level cases.
-In parallel, religion and race are under-represented, respectively in 10/12 and 8/12 of cases.
-We also look at ambiguity, considering that a hateful example is ambiguous if not all three annotators labeled it as hateful.
-The share of ambiguous content is significantly higher in false negatives compared to the whole set of hateful examples across languages and countries (t(11) = 3.76; p-value < 0.05).
+Next, we analyze false negatives, defined as hateful examples missing from the top 1% of the best model’s score distribution for each language and country.
+We first examine the targets of these tweets to compare their representation in false negatives to their overall prevalence in hate speech for each language or country.
+We find that politics and gender are overrepresented in false negatives, respectively in 10/12 and 7/12 of cases.
+Conversely, religion and race are underrepresented in 10/12 and 8/12 of cases.
+We also consider ambiguity, defining a hateful example as ambiguous if not all three annotators labeled it as hateful.
+We find that the share of ambiguous content is significantly higher in false negatives than in all hateful examples (t(11) = 3.76; $p<0.05$).
 Target academic focus and prevalence
-Motivated by the overrepresentation of certain targets in false negatives, we inspect the role of target-level academic focus, as a proxy for the distribution of hate types seen by the model during training, on performance.
-We first examine differences at the target level between the attention each target group has received in academic work and its prevalence in HateDay across languages and countries.
+Motivated by the overrepresentation of certain targets in false negatives, we inspect the role of target-level academic focus, as a proxy for the distribution of hate targets seen by the model during training, on performance.
+We first examine differences at the target level between the prevalence of each target in academic datasets and its prevalence in HateDay across languages and countries.
 We use data from a recent survey which documents the share of all hate speech datasets by target category and language Yu et al.
@@
 For each language and country, we provide a comparison between the target-level shares of (i) all hate speech datasets and of (ii) all hate speech in HateDay in Figure 6 in the Appendix.
-Our most notable finding is that political hate speech is much more prevalent in HateDay in two-thirds of cases than it is studied in academic work on hate speech.
+Our most notable finding is that political hate speech is much more prevalent in HateDay for two-thirds of the 12 languages or countries than in academic datasets.
 To a lesser extent, we also find that gender-based hate is often more prevalent in HateDay than is studied while religion-based hate is less prevalent than is studied in existing hate speech datasets.
+Target alignment and performance
 Finally, we explore the impact of the alignment between the target focus of academic work and the target prevalence in real-world social media data on performance.
 We estimate such alignment for each language and country by computing the cosine similarity between the vectors containing the share of all hate speech datasets by target focus and the share of all hate speech in HateDay by target.
-We find a strong and significant positive correlation between this alignment and detection performance (Pearson’s r: 0.76, p-value: 0.003).
+We find a strong and significant positive correlation between this alignment and detection performance (Pearson’s r: 0.76, $p=0.003$).
 This indicates that the better hate speech detection resources reflect hate target coverage in real-world social media data for a given language or country, the higher the detection performance.
-In contrast, we do not find a significant correlation between the amount of annotated datapoints in the aforementioned supersets and performance on HateDay (Pearson’s r: -0.47, p-value: 0.127).
-This indicates that the amount of detection resources plays a smaller role in cross-lingual and geographic performance differences.
+In contrast, we do not find a significant correlation between the amount of annotated datapoints in the aforementioned supersets and performance on HateDay (Pearson’s r: -0.47, $p=0.127$).
+This indicates that the amount of detection resources may play a smaller role in cross-lingual and geographic performance differences compared to the alignment between academic target focus and real-world target prevalence.
 4.3 Feasibility of Hate Speech Moderation
 Figure 4: Cost-recall tradeoff in human-in-the-loop moderation.
-Share of all HateDay tweets flagged as hateful and reviewed by moderators (%) versus share of all hate in HateDay moderated (%).
-We use the best performing model on HateDay for each language and country (Tables 2 and 3).
+Share of all HateDay tweets flagged as hateful and reviewed by moderators (%) versus share of all moderated hate in HateDay (%).
+We use the best model on HateDay for each language and country (Tables 2 and 3).
 The dashed line indicates performance of a model that flags tweets as hateful at random.
@@
 (2023) and have done so since their inception Gillespie (2018).
-In light of our performance results, we investigate the feasibility of hate speech moderation using publicly available hate speech detection models.
-Specifically, we consider human-in-the-loop moderation, which is an approach used by large social media platforms Avadhanula et al.
+In light of our poor performance results (§4.1), we investigate the feasibility of hate speech moderation using publicly available hate speech detection models.
+Given the demonstrated low performance of hate speech detection in real-world social media settings, automatic moderation, whereby content flagged as hateful by a detection model is directly moderated, may be too error-prone.
+We therefore consider human-in-the-loop moderation, which is a more realistic approach used by large social media platforms Avadhanula et al.
 (2022) whereby content likely to be hateful is flagged by detection models for review by human moderators.
-We study the trade-off between the amount of human reviewing required and the corresponding share of all hate successfully moderated (Figure 4, details in E.3).
-We find that successfully moderating a large share of all hateful content (>80%) using the best publicly available detection models would require human review of least 10% of all tweets posted in a given day.
+We study the trade-off between the amount of human reviewing required and the corresponding share of all hate successfully moderated (Figure 4, details in §E.3).
+We find that successfully moderating a large share of all hateful content (>80%) using the best publicly available detection models would require human review of at least 10% of all daily tweets.
 For such a review workload, the total share of hate successfully moderated varies between 70% and 90% for most languages or countries, apart from Turkish where this share is only 40%.
-10% of all daily tweets in a given language represents a massive amount of posts, ranging from 19.3 thousand for Kenya to 5.1 million for English.
-A smaller share of human review would only allow to moderate a small share of all hate: when flagging and reviewing 1% of all tweets posted in a given day, only 20-40% of all hate are moderated for most languages and countries.
+10% of all daily tweets may represent a massive amount of posts depending on the context and ranges from 19,300 tweets for Kenya to 5.1 million for English.
+A smaller share of human review would lead to most hate being left unmoderated: when flagging and reviewing 1% of all tweets posted in a given day, only 20-40% of all hate is moderated for most languages and countries.
 5 Discussion
@@
 (2022).
-This contrasts with the high share of social media users who report that they often encounter hate speech in their feeds Vidgen et al.
-(2019); UNESCO (2023) and points to the potential role of recommendation algorithms in amplifying harmful content Milli et al.
+This contrasts with recent survey results where two thirds of social media users report that they often encounter hate speech in their feeds UNESCO (2023) and points to the potential role of recommendation algorithms in amplifying harmful content Milli et al.
 (2025).
 Performance overestimation
-We find that traditional methods largely overestimate performance in randomly drawn samples of real-world social media data, which is very low, especially for non-European languages.
+We find that traditional evaluation of detection models, on academic datasets and functional tests, largely overestimates performance in randomly drawn samples of real-world social media data.
+Performance on real-world data is very low, especially for non-European languages.
 While past work has discussed the risk of such overestimation Arango et al.
@@
 This applies more broadly to the detection of harmful content on social media, such as misinformation Thorne et al.
+(2018); Magomere et al.
 (2025), as well as the evaluation of the bias and safety of generative large language models Ibrahim et al.
@@
 (2024); Röttger et al.
-(2024), both of which have relied so far on non-representative benchmarks for evaluation.
+(2024, 2025), both of which have relied so far on non-representative benchmarks for evaluation.
+Supervised beats zero-shot
+Despite growing enthusiasm for decoder-based large language models (LLMs), we find that they underperform supervised learning for hate speech detection—echoing findings from prior work Nozza (2021).
+This also reinforces the conclusions of a recent survey of NLP practitioners, which highlights the continued importance of annotated data for maximizing model performance Romberg et al.
+(2025).
 Dominance of Perspective API
-We find that the Perspective API often outperforms other models in hate speech detection across languages, contradicting existing evaluations on traditional hate speech datasets showing that Perspective is outperformed by academic hate speech classifiers Wich et al.
+We find that the Perspective API often outperforms other models in hate speech detection across languages, contradicting existing evaluations on traditional hate speech datasets, which showed that Perspective is outperformed by academic hate speech classifiers Wich et al.
 (2022); OFCOM (2024).
@@
 (2023).
-Yet, we also observe that Perspective’s dominance does not hold at the country level.
-This highlights the limits of the language-level approach in developing Perspective and more generally hate speech classifiers Tonneau et al.
-(2024a), which may not be tailored to national contexts where non-US English and code-mixing is prevalent.
+Still, we find that overall performance remains low, which echoes past evaluations of Perspective’s performance and biases Nogara et al.
+(2023); Hartmann et al.
+(2025).
+Also, we observe that Perspective’s dominance does not always hold at the country level, with the open-source hate speech classifier outperforming Perspective in the case of Nigeria.
+This may be due to conceptual alignment: the HateDay annotation guidelines were partly based on the same definition and instructions used to train the Nigerian classifier Tonneau et al.
+(2024b).
+This also highlights the limits of the language-level approach in developing Perspective and more generally hate speech classifiers Tonneau et al.
+(2024a), which may not be tailored to national contexts where non-US English and code-mixing is prevalent, as already demonstrated in past work Ghosh et al.
+(2021); Haber et al.
+(2023).
 The problem of offensive language
 We find that a major reason for the poor performance of hate speech classifiers on HateDay is the high prevalence of offensive content at the top of the hate score distribution, which reduces precision.
-While offensive content differs from hate speech because it does not target based on the perceived belonging of a particular characteristic, it shares semantic similarities such as the use of swear words and profanities.
+While offensive content differs from hate speech because it does not target an individual or a group based on the perceived belonging of a particular characteristic (e.g., gender, race), it shares lexical features such as the use of swear words and profanities.
 These similarities likely explain the high hate scores given by classifiers to offensive content.
 While prior work initially documented this issue Davidson et al.
-(2017), our findings demonstrate its heightened relevance in social media settings, aggravated by the substantially higher prevalence of offensive content compared to hate speech.
+(2017), our findings demonstrate its heightened relevance in real-world social media settings, as it is aggravated by the substantially higher prevalence of offensive content compared to hate speech.
 Unseen targets of hate
 Our findings reveal a misalignment between the targets of hate speech studied in academic work and their real-world prevalence in HateDay.
-Notably, political-related hate is understudied in academic work compared to its prevalence in HateDay, expanding on past work that points to a low representation of certain targets in hate speech detection resources Yu et al.
+Notably, political-related hate is understudied in academic work compared to its prevalence in HateDay, expanding past work that points to a low representation of certain targets in hate speech detection resources Yu et al.
 (2024).
@@
 In the process, political criticism and hate speech must be clearly differentiated to avoid accusations of politically biased moderation Vogels et al.
-(2020), and hate speech laws being misused for political censorship Strossen (2018).
+(2020) and hate speech laws being misused for political censorship Strossen (2018).
+Cross-geographic disparities
+Consistent with prior research Röttger et al.
+(2022a), we find that detection performance is generally lower for non-European languages, such as Arabic, Turkish and Indonesian.
+Our analysis extends this finding by offering one explanation mentioned in the past paragraph: performance disparities may stem less from varying amounts of annotated training data and more from a misalignment between the target categories prevalent in academic datasets and those that are prevalent in real-world content.
+This insight may also explain counterintuitive results such as the surprisingly low real-world performance for the English and US contexts and higher performance for India, despite English and the US having the most annotated resources Tonneau et al.
+(2024a).
+In real-world English-speaking and US contexts, politically motivated hate is among the most prevalent forms of hate speech; yet, it is underrepresented in US-dominated English academic datasets.
+By contrast, in the Indian context, there is greater alignment between the types of hate emphasized in academic datasets and those commonly encountered in practice, which may help explain the observed performance differences.
+Furthermore, the overall prevalence of hate speech is significantly lower in English and US data (0.4% and 0.5% of posts, respectively) compared to Indian data (1.7%), suggesting that our estimate of model performance in English and US settings may be less robust.
 Other challenges for hate speech detection
-Our analysis identifies additional challenges for hate speech detection partly identified in past work: First, the difficulty of distinguishing between the use and the mention of hate speech, for instance as part of counter speech Röttger et al.
+We present further challenges for hate speech detection that have partly been identified in past work: First, the difficulty of distinguishing between the use and the mention of hate speech, for instance when slurs or hateful terms are quoted as part of counter speech or educational content Röttger et al.
 (2021); Gligoric et al.
@@
 (2024).
-Second, the inability of classifiers to capture subtle differences in wording that change a statement from hateful to non-hateful.
-Third, the lack of context in many posts, especially replies, poses challenges to classification, and calls for more contextualized hate speech detection Pérez et al.
+Second, the inability of classifiers to capture subtle differences in wording and phrasing that can make a non-hateful statement become hateful Sap et al.
 (2019).
+Third, the lack of context in many posts—especially replies—poses challenges to classification and calls for more contextualized hate speech detection Pérez et al.
+(2023).
+Our analysis underlines the relevance of these challenges, given their prevalence in real-world settings, and highlights the need to address them as a priority.
 Implications for hate speech moderation
-Our results indicate that hate speech moderation using public hate speech detection models is ineffective at curbing online hate speech at a limited cost.
-Indeed, low real-world detection performance renders fully automated moderation ineffective and potentially harmful, as it fails to protect users from hate speech while likely censoring non-hateful content, such as counter-speech, confirming concerns from past theoretical work Gorwa et al.
+Our results indicate that fully automated hate speech moderation using public hate speech detection models is undesirable.
+Indeed, low real-world detection performance renders fully automated moderation ineffective and potentially harmful, as it fails to protect users from hate speech while likely removing non-hateful content, such as counter-speech.
+This reinforces concerns raised in earlier theoretical work Gorwa et al.
+(2020) and complements past empirical findings highlighting the limitations of hate speech detection research for content moderation Ye et al.
+(2023); Zheng et al.
 (2024).
-We also show that human-in-the-loop moderation based on public detection models can only effectively moderate a large share of hate speech across our languages and countries of interest at a prohibitively high annotation cost, extending past results for the Nigerian Twitter context Tonneau et al.
+We also show that human-in-the-loop moderation based on public detection models can effectively moderate a large share of hate speech across our languages and countries of interest but that this implies having humans review a non-trivial share of all daily content, extending past results for the Nigerian Twitter context Tonneau et al.
 (2024b) to more languages and countries.
-Beyond financial costs, the required scale of human review also raises concerns about potential reviewer harm from repeated exposure to harmful content Kirk et al.
+Beyond financial costs, the required scale of human review also raises concerns about potential reviewer harm from repeated exposure to harmful content Roberts (2016); Kirk et al.
 (2022).
-Our results also support the claim that detection alone will not solve the hate speech problem Parker and Ruths (2023) and call for complementary solutions to this problem.
-These include preventive approaches which aim to reshape speech norms and have proven effective to curb hate speech, for instance by prompting users to reconsider harmful posts Katsaros et al.
-(2022), or by confronting hate speech with counter speech Munger (2017); Hangartner et al.
+Our results support the claim that detection alone will not solve the hate speech problem Parker and Ruths (2023) and call for complementary solutions.
+These include preventive approaches that aim to reshape speech norms and have proven effective in curbing hate speech, for instance by prompting users to reconsider harmful posts Katsaros et al.
+(2022) or by confronting hate speech with counter speech Munger (2017); Hangartner et al.
 (2021).
-Finally, we note that our results do not imply that hate speech moderation itself is ineffective, because social media platforms use their own private models for moderation and have access to more data which may improve performance (see Limitations).
-Yet, without access to real-world performance metrics for platform models, the effectiveness of hate speech moderation on social media remains uncertain.
 6 Conclusion
 In this article, we introduced HateDay, the first global hate speech dataset representative of real-world social media settings.
-We found that evaluations based on previous academic datasets overestimate real-world performance, which is very low.
-We further discussed the implications of our results for hate speech moderation, namely that it is unfeasible using public detection models.
-In light of our results, future work is needed to improve hate speech detection performance.
-In the process, we encourage researchers to ground their evaluation in the real-world settings in which their models may eventually be used.
-We also encourage platforms to be more transparent on how their detection models perform in social media contexts, in order to better understand the effectiveness of hate speech moderation.
-We hope that our dataset and experimental results will stimulate work in both of these directions to help solve this important problem.
+Using HateDay, we showed that evaluating hate speech detection models on standard academic datasets substantially overestimates their real-world performance, which is very low.
+We explored the implications of this finding for content moderation, concluding that relying on public detection models for automated moderation is currently ill-advised due to their high error rate.
+Accordingly, improving real-world model performance should be a key focus of future research.
+We also found that human-in-the-loop moderation can be more accurate, but only if a substantial portion of daily content is manually reviewed—raising important questions about the feasibility and desirability of such an approach at scale, which merits further investigation.
+Ultimately, we urge researchers to evaluate future detection models within the real-world contexts where they are likely to be deployed.
+We also call on platforms to provide greater transparency about how their detection systems perform in real social media environments, to better assess moderation effectiveness.
+We hope that our dataset and findings will drive progress in both areas and contribute to addressing this pressing challenge.
 Limitations
@@
 Low number of positives: The random samples in HateDay used to evaluate hate speech detection in real-world settings contains a low number of hateful examples, ranging from 31 for Kenya to 430 for India.
-This low number is linked to the very low prevalence of hate speech in our dataset as well as our budget constraint which impeded us from further expanding the annotation effort.
+This low number is linked to the very low prevalence of hate speech in our dataset as well as our budget constraint, which impeded us from further expanding the annotation effort.
 While statistically significant, we acknowledge that our performance results on HateDay are necessarily more uncertain, as illustrated by the larger confidence intervals in Tables 5 and 6.
@@
 This limits the generalizability of our performance results to data from other social media platforms and covering other timespans.
-Online hate speech is multimodal: Our work focuses only on text-based hate to limit annotation costs, but we acknowledge that online hate speech is multimodal and that a non-trivial share of this phenomenon, which our analysis necessarily misses, is expressed through images, sounds, or videos Hee et al.
+Online hate speech is multimodal: Our work focuses only on text-based hate to limit annotation costs, but we acknowledge that online hate speech is multimodal and that a non-trivial share of this phenomenon, which our analysis necessarily misses, is expressed through images, sounds, or videos Botelho et al.
+(2021); Hee et al.
 (2024).
@@
 This implies that the representativeness of HateDay may not be perfect, with language or country sets containing some posts in other languages or from other countries.
-We acknowledge this limitation but argue that HateDay remains as close as a set of representative samples of Twitter as can be with existing stratification tools.
+We acknowledge this limitation, but argue that HateDay remains the most representative samples of Twitter possible with current stratification tools.
 Moderation prior to data collection: Our analysis assumes that the hateful content in HateDay is representative of all hateful content posted on Twitter on the day of analysis.
@@
 (2023), we believe that the enforcement of moderation in such a short timeframe is likely to be minimal.
+Experiments
+Other prompts could lead to different results: We craft a prompt using the terms “hateful” or “non-hateful” (see D.2 for details), which exhibit good performance in past research for hate speech detection using zero-shot learning Plaza-del arco et al.
+(2023).
+We do not test other prompts and acknowledge that using other prompts may have an impact on classification performance.
 Moderation analysis
@@
 While we consider only public hate speech detection algorithms as a way to flag hate speech in the flow of social media content, we are also aware that platforms use other flagging mechanisms based on private data, such as user reporting or banned word lists, which may increase the share of all hate moderated for a given annotation cost.
-Finally, the metric platforms aim to reduce is the view count of harmful content such as hate speech, which is arguably more relevant than the sheer prevalence of it from a harm perspective but is also private for the considered timeframe.
+Finally, the metric platforms aim to reduce is the view count of harmful content such as hate speech, which is arguably more relevant than the sheer prevalence of it from a harm perspective.
+Unfortunately, such view data is unavailable to us for the considered timeframe.
 Ethical Considerations
@@
 Intended use
-The HateDay dataset will be shared publicly upon publication and its intended use is for research purposes only.
+The intended use of the HateDay dataset is for research purposes only.
 Acknowledgements
 We thank the annotators for their excellent research assistance.
-We also thank Dhruti Sanghavi for outstanding visual design work.
-The study was supported by funding from the United Kingdom’s Foreign Commonwealth and Development Office (FCDO), the World Bank’s Research Support Budget, and the Bill and Melinda Gates Foundation.
+The study was supported by funding from the United Kingdom’s Foreign Commonwealth and Development Office (FCDO), the World Bank’s Research Support Budget, and the Gates Foundation (INV057844).
 This work was also supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
-The findings, interpretations, and conclusions expressed in this paper are entirely those of the authors.
+The findings, interpretations, and conclusions expressed in this article are entirely those of the authors.
 They do not necessarily represent the views of the International Bank for Reconstruction and Development/World Bank and its affiliated organizations or those of the Executive Directors of the World Bank or the governments they represent.
+Scott Hale was supported by the ESRC Digital Good Network (grant reference ES/X502352/1).
+Paul Röttger is a member of the Data and Marketing Insights research unit of the Bocconi Institute for Data Science and Analysis, and is supported by a MUR FARE 2020 initiative under grant agreement Prot.
+R20YSMBZ8S (INDOMITA).
 Appendix A Twitter Day
@@
 B.2 Annotation guidelines
+We provide a version of the annotation guidelines at https://github.com/manueltonneau/hateday_acl/blob/main/annotation_guidelines.pdf
 B.3 Recruitment and compensation
@@
 The language-level supersets are documented in Tonneau et al.
-(2024a) and obtained from Hugging Face [^3] https://huggingface.co/collections/manueltonneau/hate-speech-supersets-664ef6d2bc40cce7a8b1092f.
-The number of annotated datapoints in these supersets for each language is:
+(2024a) and obtained from Hugging Face.
+[^2] https://huggingface.co/collections/manueltonneau/hate-speech-supersets-664ef6d2bc40cce7a8b1092f The number of annotated datapoints in these supersets for each language is:
   • Arabic: 464,260
@@
 First, we inspect the Hate Speech Data Catalogue [^3] https://hatespeechdata.com/ Vidgen and Derczynski (2020).
-Second, we inspect the datasets listed in the latest survey of HS datasets Poletto et al.
+Second, we inspect the datasets listed in the latest survey of hate speech datasets Poletto et al.
 (2021).
@@
   3. 3.
-    The dataset focuses on HS, defined broadly as “any kind of communication in speech, writing or behavior, that attacks or uses pejorative or discriminatory language with reference to a person or a group on the basis of who they are, in other words, based on their religion, ethnicity, nationality, race, color, descent, gender or other identity factor” UN (2019).
+    The dataset focuses on hate speech, defined broadly as “any kind of communication in speech, writing or behavior, that attacks or uses pejorative or discriminatory language with reference to a person or a group on the basis of who they are, in other words, based on their religion, ethnicity, nationality, race, color, descent, gender or other identity factor” UN (2019).
 C.2.2 Target categorization
@@
 D.4 Number of parameters
-The supervised learning models are largely based on BERT base which has 110 million parameters. The decoder-based models (Aya and Llama3.1) both have 8 billion parameters.
+The supervised learning models are largely based on the BERT base architecture which has 110 million parameters. The decoder-based models (Aya and Llama3.1) both have 8 billion parameters.
 Appendix E Additional Results
@@
 Arabic IbrahimAmin/marbertv2-finetuned-
-egyptian-hate-speech-detection 25.8±0.6 76.9 $\pm$2.0 5.8$\pm$3.0
+egyptian-hate-speech-detection 25.8±0.6 76.9
+$\pm$2.0
+5.8
+$\pm$3.0
 Hate-speech-CNERG-
@@
 Figure 7: Offensive-to-hateful count ratio versus the share of offensive content in top tweets (in %).
-Top tweets are defined as the 1% (N=200) top-scored tweets in HateDay using the best performing model on the latter.
+Top tweets are defined as the 1% (N=200) top-scored tweets in HateDay using the best performing model on the dataset.